Astrid1:

@sNoWfLaKe0531 ss below e.e I need help with both of them...

Astrid1:

snowflake0531:

n and p? what? where?...

snowflake0531:

Let's do number 4 xd

Astrid1:

Alright..

snowflake0531:

So you know that for example, sqrt10... you know that 3 squared = 9, and 4 squared equals 16, so sqrt10 is between 3 and 4

Astrid1:

Ok...

snowflake0531:

So then for sqrt54 which two numbers squared? Go through it like \[1^2 = 1\] \[2^2 =4\]\[3^2=9\]\[4^2=16\]\[5^2=25\]\[6^2=36\]\[7^2=49\]\[8^2=64\]\[9^2=81\]\[10^2=100\]

Astrid1:

\[7^2~and~8^4?\]

Astrid1:

*8^2

snowflake0531:

yep, 7 and 8

Astrid1:

Ok..

snowflake0531:

So then do the same for all of htem xd

Astrid1:

Alright, Thx ilysm (no weirdness) ur so smort and helpful.

snowflake0531:

lolll, thxx and yw~